Financial position 2016: Machinery = $50,600

Some machinery (a non-current asset), which was bought on 1 January 2013 for $26,000, has proved to be unsatisfactory. It was part-exchanged for some new machinery on 1 January 2014, and WW Associates paid a cash amount of $12,000. The new machinery would have cost $30,000 had the business bought it without the trade-in. The business uses the reducing-balance method of depreciation for non-current assets at the rate of 30% each year

How to calculate machinery depreciation step by step?
Answer #1

Depreciation in 2013 = 26000*30% =7800

Depreciation in 2014 =30000*30% =9000

Depreciation in 2015 =30000-9000 *30% =6300

Financial position of machinery is 50600 in 2016,hence depreciation =50600*30% =15180